Quadrilateral abcd is a rhombus with an area of 600 square inches and diagonal ac=30 what is the perimeter of quadrilateral abcd?

Answer by zeti68(10)

Hi,
(ac)(bd)/2=600
ac=30
30(bd)=1200
bd=40
Measure of the diagonals are 30 and 40
30/2=15 (half of the diagonal ac)
40/2=20 (half of the diagonal bd)
side = sq(15^2+20^2)= sq(625)=25
perimeter = 4(25)=100
Answer: perimeter = 100 inches